P3(25pts). If f(x,y)=x2 +4x -y3 -xy2 ,
determine all critical vectors of f and using the second derivative test classify all local extrema.
Answer #1

f(x,y) = x² + 4x -43 _ xy2 Differentiating partially with respect to x64 We get, f = 2x+4-y2 fy = - 3y*=2XY fax=2, fyy = -64NW at (2); fare for – Farm * = 2,6-6-(1) - 2+(-3) -(-2:00) .. . = 2(+6+3) - 4 = -6-4.50 : --(-3,1) is also a saddle point